OpenStudy (anonymous):

Which of these is equal to cos 47°? A) cos 43° B) sin 43° C) tan 43° D) cos 133°

OpenStudy (hoblos):

cos(a) = sin( 90-a)

OpenStudy (hoblos):

so cos 47° = sin (90-47) = sin 43°

OpenStudy (anonymous):

hah were you taking a usatest prep quiz